the sum of the solutions of 3x^+3x=36
Answer by Nate(3500)

3x^2 + 3x = 36
x^2 + x = 12
(x + 0.5)^2 = 12.25
x = -0.5 +- 3.5
x = -4 and x = 3
Sum ~> -4 + 3 = -1